OpenStudy (anonymous):

Determine if the Mean Value Theorem applies, then find "c" such that f'(c)=(f(b)-f(a))/(b-a)

OpenStudy (anonymous):

f(x)=x^3-x^2-2x for [-1,1]

OpenStudy (anonymous):

why would the answer be c=-1/3 and not 1 @TweT226

OpenStudy (anonymous):

you have to take the derivative

OpenStudy (anonymous):

\[f'(c)=\frac{ f(b)-f(a) }{ b-a }\] \[3c^2-2c-2=\frac{ -2-0 }{ 1- (-1)}\]

OpenStudy (anonymous):

\[3c^2-2c-2=-1\] \[3c^2-2c-1=0\] \[(3c+1)(c-1)=0 \] \[c=\frac{ -1 }{ 3 }, 1\]

OpenStudy (anonymous):

so how come my teacher chose \[\frac{ -1 }{ 3}\] over c=1?

OpenStudy (anonymous):

I think it's because it's right on the endpoints

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

i mean interval [-1, 1]

OpenStudy (anonymous):

the theorem states that the equation is differentiable between (a,b) not [a,b] the domain is [=1,1] in the problem, so when you take the derivative, c cannot equal a or b.

OpenStudy (anonymous):

oh u mean it's in between

OpenStudy (anonymous):

oh ok gotcha ;)

OpenStudy (anonymous):

-1*

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@yun2thejae can u help me with another problem pls? ^^

OpenStudy (anonymous):

sure, if you would like.